HIGH RESOLUTION MICROWAVE SPECTROSCOPY OF THE ALLENE DERIVATIVES H_{2}C_{3}HC_{2}H, H_{2}C_{3}HC_{4}H, AND H_{2}C_{3}HC_{3}N

\underline{HARSHAL GUPTA}, Division of Engineering and Applied Sciences, Harvard University, Cambridge, Massachusetts 02138 and Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, The University of Texas at Austin, Austin, Texas 78712; M. C. McCARTHY and P. THAD-DEUS, Harvard-Smithsonian Center for Astrophysics, Cambridge, Massachusetts 02138 and Division of Engineering and Applied Sciences, Harvard University, Cambridge, Massachusetts 02138.

Molecules of the form H_{2}C_{3}H–R (where R is a carbon chain) are of astrophysical interest in light of the recent centimeterwave band detection of cyanoallene (H_{2}C_{3}HCN) towards TMC-1^{a}. The carbon chain molecules H_{2}C_{3}HC_{2}H, H_{2}C_{3}HC_{4}H, and H_{2}C_{3}HC_{3}N have been investigated between 5 and 41 GHz by Fourier transform microwave spectroscopy of a supersonic molecular beam. Accurate rotational and centrifugal distortion constants have been derived for all three molecules from their a- and b-type transitions, and owing to the high spectral resolution of this technique, nitrogen hyperfine structure has been resolved for H_{2}C_{3}HC_{3}N. Several ^{13}C isotopic species have now been observed in natural abundance, suggesting that accurate determinations of experimental structures may be feasible. Because these molecules are structurally similar to known astronomical molecules and possess large dipole moments, they are good candidates for astronomical detection.

aF.J. Lovas, Anthony J. Remijan, J.M. Hollis, P.R. Jewell, and L.E. Snyder ApJ, \underline{637}: L37-L40, (2006).